OpenStudy (anonymous):

5x+2y=-3 3x+3y=9 what's the value of X and Y?

OpenStudy (anonymous):

step1; multiply first equation with 3 step2; multiply second equation with -2 step3; add the two equations see what you get

OpenStudy (anonymous):

i don't get it when you say multiply them together? Isn't it add or subtract the opposites?

OpenStudy (anonymous):

i never said multiply the equations

OpenStudy (anonymous):

alright. when you said multiply first equation with 3, though can you explain what you mean?

OpenStudy (anonymous):

you'll get 15x + 6y = -9

OpenStudy (anonymous):

so, then subtract 15x from -9?

OpenStudy (anonymous):

or would i subtract 6?

jimthompson5910 (jim_thompson5910):

when he said "multiply the first equation by 3", he means 5x+2y=-3 3(5x+2y)=3(-3 ) ... Multiply both sides by 3. 15x + 6y = -9 ... distribute

jimthompson5910 (jim_thompson5910):

So 5x+2y=-3 turns into 15x + 6y = -9

OpenStudy (anonymous):

Ohhhh, i got it now (: thank you so much for clearing that up!

jimthompson5910 (jim_thompson5910):

Now multiply both sides of the second equation by -2 3x+3y=9 -2(3x+3y)=-2*9 -6x - 6y = -18

jimthompson5910 (jim_thompson5910):

This means that 5x+2y=-3 3x+3y=9 becomes 15x + 6y = -9 -6x - 6y = -18 The whole reason we did all this was to add up the equations so one variable cancels out. Notice how 6y + (-6y) = 0y and that goes away leaving you to be able to solve for x

OpenStudy (anonymous):

so x=-9 and y=-18?

jimthompson5910 (jim_thompson5910):

no that's not correct

jimthompson5910 (jim_thompson5910):

Add the two equations 15x + 6y = -9 -6x - 6y = -18 ---------------- 9x + 0y = -27 9x = -27 x = ???

OpenStudy (anonymous):

x=3

jimthompson5910 (jim_thompson5910):

closer, but still no

OpenStudy (anonymous):

ohh, -3 ahaha

jimthompson5910 (jim_thompson5910):

better, now use that to find y

OpenStudy (anonymous):

so plug in -3 to 2y=-3? making it -6=-3 dividing that and making it 2?

jimthompson5910 (jim_thompson5910):

pick any equation where there's an x and y so I'm picking the first one 5x+2y=-3 5(-3)+2y=-3 .. plug in x = -3 -15+2y=-3 2y = -3 + 15 2y = 12 y = 12/2 y = 6 So the solutions are x = -3 and y = 6

jimthompson5910 (jim_thompson5910):

They form the ordered pair (-3, 6) Basically, the two lines cross at the point (-3, 6)

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome